OpenStudy (anonymous):

I will give you a medal if can help me Solve the system using the elimination method. -2x+5/3y=-5 3x=7/2y=9/2 Show work PLEASE :)

OpenStudy (dumbcow):

have you tried this on your own? what did you get

OpenStudy (anonymous):

get rid of the annoying fractions to make life easier. multiply first one by 3, second by 2. then see what you get

OpenStudy (anonymous):

2nd eqution looks incorrect

OpenStudy (anonymous):

yes sorry it is it's 3x-7/2=9/2

OpenStudy (anonymous):

\[-6x+5y=-15\] \[6x-7y=9\]

OpenStudy (anonymous):

add and the x's drop out, giving yu \[-2y=-6\] sp \[y=-3\]

OpenStudy (anonymous):

then replace y by -3 in either equation (the ones without the fractions would be easiest) so solve for x.